# https://kartik-3004.github.io/facexformer_web/

### v1

@@ 4.1 Datasets and Metrics
We perform co-training, where the model is simultaneously trained for multiple tasks using a total of ten datasets with task-specific annotations. We conduct intra-dataset and cross-dataset evaluations and present our results on the test sets according to the standard protocol for each task using the following datasets:
Training: Face Farsing: LaPa [41], CelebAMaskHQ [34] Landmarks Detection: 300W [53], WFLW [76]; Head Pose Estimation: 300W-LP [93], AFLW2000 [94], Attributes Recognition: CelebA [42]; Age/Gender/Race estimation: UTKFace [85], FairFace [28]; Landmarks Visibility Prediction: COFW [5].
Test (Intra-dataset): Face Parsing: LaPa, CelebAMaskHQ; Landmarks Detection: 300W, WFLW; Attributes Recognition: CelebA; Age/Gender/Race Estimation: UTKFace, FairFace; Landmarks Visibility Prediction: COFW.
Test (Cross-dataset): Landmarks Detection: 300VW [57]; Head Pose Estimation: BIWI [16]; Attributes Recognition: LFWA [74]; Age/Gender Estimation: FFHQ [29]. The evaluation metrics used are F1-score for face parsing, Normalized Mean Error (NME), Failure Rate (FR@$10\%$) and Area Under the Curve (AUC@$10\%$) for landmarks detection, Mean Absolute Error (MAE) for head pose estimation, and accuracy for attributes recognition, age/gender/race estimation and landmarks visibility prediction.
@@ 4.2 Implementation Details
We train our models using a distributed PyTorch setup on eight A5000 GPUs, each equipped with $24$GB of memory. The models’ backbones are initialized with ImageNet pre-trained weights and processes input images at a resolution of $224\times 224$. We employ the AdamW optimizer with a weight decay of $1e^{-5}$. All models are trained for $12$ epochs with a batch size of $48$ on each GPU, and an initial learning rate of $1e^{-4}$, which decays by a factor of $10$ at the $6^{th}$ and $10^{th}$ epochs. For data augmentation, we randomly apply Gaussian blur, grayscale conversion, gamma correction, occlusion, horizontal flipping, and affine transformations, such as rotation, translation and scaling. The number of FaceX decoder $N$ is set to two. To ensure stable training across tasks when using multiple datasets of varying sample sizes, we equalize the representation of each task’s samples in every batch through upsampling. Additional details on our implementation are provided in the supplementary document.
@@ 4.3 Main results
Methods LaPa CelebAMaskHQ 300W WFLW BIWI CelebA FairFace UTKFace COFW
F1-score F1-score NME $\text{FR}_{@10\%}$ MAE Acc. (Age)[Gen]{Race}Acc. (Age)[Gen]{Race}Acc. Acc.
$\text{EAGR}_{473}$[65] 91.07 90.84 - - - - - - -
$\text{AGRNET}_{473}$ [64] 92.34 90.67 - - - - - - -
[…]
Wing [17] - - - 6.00 - - - - -
SBR [14] - - 4.10 - - - - - -
DeCaFa [12] - - 3.39 4.84 - - - - -
HRNet [69] - - 3.32 4.64 - - - - -
STAR Loss [90] - - 2.87 2.32 - - - - -
FDN [83] - - - - 3.93 - - - -
WHENet [89] - - - - 3.81 - - - -
TriNet [7] - - - - 3.97 - - - -
img2pose [2] - - - - 3.78 - - - -
6DRepNet [22] - - - - 3.47 - - - -
TokenHPE [82] - - - - 3.72 - - - -
SSP+SSG [26] - - - - - 88.24 - - -
SSPL [58] - - - - - 91.77 - -
FairFace [27] - - - - - - (59.7)/[94.2]/{78.01} (61.7)/[92.5]/{83.9} -
MiVOLO [32] - - - - - - (61.07)/[95.73] - -
SoftCon [23] - - - - - - - {87.00} -
EnD [63] - - - - - - - {88.40} -
ProS [13] 92.00 - 3.27 - - 91.88 - - -
FaRL [88] 91.62 90.34 3.35 5.72 - 91.39 (61.78)/[95.65] - -
HyperFace [51] - - 3.62 6.09 3.46 - [94.26] [95.52] 88.91
AllinOne [52] - - 4.52 15.45 5.64 - (54.18)/[88.22] (61.48)/[93.70] 84.11
FaceXFormer 91.10 89.83 3.53 4.56 3.65 91.79 (59.38)/[95.20]/{77.91} (63.93)/[95.69]/{87.45} 90.16
Table 2: Comparison with specialized models and existing multi-task networks
In Table 2, we present a comparative analysis of FaceXFormer with recent methods across a variety of tasks. A significant highlight of our work is its unique capability to deliver substantial results in all scenarios using a single unified model. FaceXFormer achieves state-of-the-art performance in face parsing (@224×224@$\times 224$ 224) and attributes recognition with a mean F1 of $90.46$ over LaPa and CelebAMaskHQ, and a mean accuracy of $91.79$ on CelebA. We observe that the performance on age/gender/race estimation is on par with some specialized bias-mitigation methods, showing that it exhibits minimal bias on these factors, even when trained on a small number of data points. This aspect is further explored in the ethical considerations section. FaceXFormer demonstrates competitive performance at par with leading methods like STAR Loss [90] and TokenHPE [82] for landmarks detection and head pose estimation respectively. The performance discrepancy in landmarks detection is mainly because we don’t employ additional heatmaps during training, which is the recent trend in landmarks detection literature. Similarly, current works utilize different head pose representations such as rotation matrix and 6D representation to improve performance. We avoid using auxiliary information and advanced representations as our goal is not to specialize in a single task, but to learn a unified representation suitable for a wide spectrum of tasks.
We also compare our work with recent models that aim at unification. ProS [13] and FaRL [88] are pretraining methods for face tasks aimed at learning a strong encoder that generates generalized feature representations. This work is tangential to HyperFace [52] and AllinOne [52], which are more similar to our proposal, i.e., learning of a unified decoder. We observe that FaceXFormer outperforms other unified model baselines across multiple benchmarks. However, HyperFace performs better in head pose estimation. We observe that different tasks complement each other and boost each other’s performance, as can be seen in the case of HyperFace. HyperFace is jointly trained for landmarks detection and head pose estimation, both of which are regression-based tasks that mutually enhance their performance. We observe a similar performance increase when FaceXFormer is trained exclusively on regression-based tasks. This enhancement is primarily because the model has learned regression-specific feature representations, which are optimized for these tasks. However, when these tasks are combined with others that do not share the same feature requirements, we notice a decrement in performance. We further explore the task synergy of FaceXFormer in the ablation section of our work.
In our work, we simultaneously train for eight heterogeneous tasks, presenting a formidable challenge compared to previous works. This challenge arises primarily because these tasks require task-specific features and pre-processing steps, given the tasks are so distinct and diverse. This ability to effectively handle a multitude of facial analysis tasks, despite the small performance gaps noted, reinforces our assertion that FaceXFormer is a unified versatile model suitable for diverse facial analysis tasks.
4.4 Qualitative results
Figure 4: Qualitative comparison of FaceXformer against other multi-task models
In Figure 4, we present the qualitative results of FaceXFormer for a selected subset of tasks and compare it against other baselines. Specifically, in the second row, we highlight our superior performance in landmarks detection. In the first row, we display the face parsing results, demonstrating tasks beyond what HyperFace and AllinOne can perform. As discussed in Section 3, FaceXFormer can be prompted with task-specific tokens during inference to obtain multiple predictions. We prompt two random in-the-wild faces for face parsing, head pose, age, gender and race estimation and show the results in Figure 3. This demonstrates FaceXFormer’s robust performance in in-the-wild scenarios.
4.5 Ablation Study
Table 3: Ablation analysis for multi-task synergy between segmentation, regression and classification
Table 4: Performance analysis with different backbones and different backbone size
Seg Reg Cls Seg Reg Cls
F1 NME Acc.
89.64 - -
- 3.84 -
- - 82.53
90.46 3.99 82.68
Backbone Seg Reg Cls
F1 NME Acc.
ResNet101 89.63 4.16 82.15
ConvNext 90.77 3.92 82.60
Swin/T 89.91 4.1 82.43
Swin/S 90.18 3.90 82.38
Swin/B 90.46 3.99 82.68
Table 4: Performance analysis with different backbones and different backbone size
In this section, we discuss how different tasks synergize to boost performance and explore the impact of various backbones and their sizes on performance. Additionally, we highlight the importance of multi-scale feature fusion.
4.6 Task Synergy
We train FaceXFormer on a subset of tasks, categorizing them into three categories: Segmentation (Seg), Regression (Reg), and Classification (Cls). Segmentation includes face parsing; regression encompasses landmarks detection and head pose estimation; and classification covers attributes recognition, age/gender/race estimation, and landmarks visibility prediction. From Table 4, we can see that segmentation and classification performance is better when trained jointly with other tasks. However, we observe a small drop in performance when tasks are trained jointly with the regression category. We attribute this to the fact that different tasks require different kinds of representations for prediction. Segmentation demands fine-grained feature representations; attributes recognition and other classification tasks require semantic information, while regression-based tasks necessitate a blend of high-level and finer features. Our proposed FaceXFormer addresses this challenge by using fused feature representations and treating multiple tasks as tokens, thereby making joint training across multiple tasks effective.
4.7 Effect of different backbone and different size for FaceXformer
In Table 4, we conducted experiments to analyze the effect of different backbones and their sizes on FaceXformer. ConvNext leads in segmentation and regression with an F1 score of 90.77% and an NME of 3.92, but is closely followed by Swin Transformer variants, which exhibit competitive performance and superior scalability. Conversely, ResNet101 shows the lowest metrics, such as an F1 score of 89.63% and an NME of 4.16, indicating its limitations in handling larger, more complex datasets and its smaller receptive field compared to Swin Transformers. The choice of Swin Transformer over ConvNext for FaceXformer is based on its enhanced scaling capabilities and superior global contextual understanding, which are crucial for facial analysis tasks, despite its slightly lower scores in certain metrics. This decision underscores the importance of fitting the backbone to the model’s architecture to improve overall task synergy and performance.
5 Discussions
In this section, we discuss the broader impact of our work. Additionally, we talk about the ethical considerations and limitations of FaceXFormer.
5.1 Broader Impact
The world is moving towards transformers because of its potential to model large amounts of data [4, 30, 3]. Presently, the face community lacks large-scale annotated datasets to train foundational models capable of performing a wide spectrum of facial tasks. The largest clean dataset, WebFace42M [96], lacks annotations for face parsing, landmarks detection and attributes recognition. FaceXFormer can be used as an annotator for large-scale data, and can be continually improved through successive rounds of annotation and fine-tuning. We aim to propel the face community towards developing foundation models that cater to a variety of facial tasks. Additionally, FaceXFormer is a lightweight model that provides real-time output based on task-specific queries and can be appended with existing facial detection systems to provide additional information. It can also serve as a valuable tool in surveillance, and provide auxiliary information for subject analysis and image retrieval.
5.2 Ethical considerations
In our work, we utilize a total of $14$ datasets for training and evaluation. We obtain these datasets following the procedures stated on their respective pages and signed the license agreements if and when necessary. As we train our models on multiple datasets designed for different tasks, the representation of subjects across different age groups, genders, and races is not equal. This imbalance may introduce bias in the model. Therefore, we provide an analysis using the FairFace [27] dataset, which is balanced in terms of age, gender and race. We follow [49] and define the "Non-white" group to include multiple racial categories: "Black", "Indian", "East Asian", "Southeast Asian", "Middle Eastern" and "Latino". As can be seen from Table 5, FaceXFormer shows the smallest performance discrepancy across different racial groups and exhibits minimal bias compared to other models, despite being trained on fewer data points. This can be attributed to race estimation being of the tasks in co-training.
@@ 5.2 Ethical considerations
Table 5: Age and gender accuracy w.r.t two race groups on FairFace
5.3 Limitations
We acknowledge the shortcomings of the proposed FaceXFormer in not achieving state-of-the-art performance in certain tasks, such as landmarks detection and head pose estimation. This is primarily because the SOTA models for these tasks utilize auxiliary heat-map based loss and advanced representations, respectively. Nevertheless, FaceXFormer uniquely handles up to eight different heterogeneous tasks within a single model and offers competitive performance across these tasks in real-time, enhancing its suitability for deployment. Additionally, while it supports tokens for various tasks, it is not yet interactive and lacks full promptability. Our future work will focus on developing an interactive foundational model capable of handling multiple tasks.
@@ 6 Conclusion
In conclusion, the FaceXformer introduces a novel end-to-end unified model that efficiently handles a comprehensive range of facial analysis tasks. By adopting a transformer-based encoder-decoder architecture and treating each task as a learnable token, our approach successfully integrates multiple tasks within a single framework. The proposed parameter-efficient decoder, FaceX, enhances the model’s ability to learn robust and generalized face representations across diverse tasks. Our comprehensive experiments show how segmentation, regression, and classification facial analysis tasks improve each other. Additionally, training on multiple datasets leads to better representation learning. In conclusion, we demonstrate that facial tasks can be treated as tokens, leading to the unification of tasks; following this, we hope our work provides a foundation for developing models capable of performing multiple facial analysis tasks.
Acknowledgements

### v2

@@ 4.1 Datasets and Metrics
We perform co-training, where the model is simultaneously trained for multiple tasks using a total of 9 datasets with task-specific annotations. We conduct intra-dataset and cross-dataset evaluations and present our results on the test sets according to the standard protocol for each task using the following datasets:
Training: Face Farsing: CelebAMaskHQ [38]; Landmarks Detection: 300W [75]; Head Pose Estimation: 300W-LP [128]; Attributes Prediction: CelebA [50]; Facial Expression Recognition: RAF-DB [42], AffectNet [58]; Age/Gender/Race estimation: UTKFace [120], FairFace [32]; Visibility Prediction: COFW [7].
Test (Intra-dataset): Face Parsing: CelebAMaskHQ [38]; Landmarks Detection: 300W [128]; Attributes Prediction: CelebA [50]; Facial Expression Recognition: RAF-DB [42]; Age/Gender/Race Estimation: UTKFace [120], FairFace [32]; Visibility Prediction: COFW [7].
Test (Cross-dataset): Landmarks Detection: 300VW [78]; Head Pose Estimation: BIWI [18]; Attributes Prediction: LFWA [101].
The evaluation metrics used are the F1-score for face parsing, Normalized Mean Error (NME) for landmark prediction, Mean Absolute Error (MAE) for head pose estimation and age estimation, accuracy for facial expression recognition, attributes prediction, gender estimation, race estimation, and recall at $80$% precision for visibility prediction.
Methods Expression Methods Visibility Methods Age (MAE)
(RAF-DB) (COFW) UTKFace
DLP-CNN [42] 80.89 RCPR [7] 40 OR-CNN [63] 5.74
gACNN [45] 85.07 Wu et al. [105] 44.43 Axel Berg et al. [4] 4.55
IPA2LT [114] 86.77 Wu et al. [104] 49.11 CORAL [8] 5.47
RAN [95] 86.90 ECT [116] 63.4 Gustafsson et al. [23] 4.65
CovPool [2] 87.00 3DDE [89] 63.89 R50-SORD [65] 4.36
SCN [94] 87.03 MNN [90] 72.12 VOLO-D1 [36] 4.23
DACL [19] 87.78 DLDL-v2 [21] 4.42
KTN [40] 88.07 MWR [79] 4.37
DMUE [77] 88.76
SwinFace [66] 86.54 SwinFace [66] $\times$ SwinFace [66] $\times$
QFace [83] 92.86 QFace [83] $\times$ QFace [83] $\times$
Faceptor [67] 87.58 Faceptor [67] $\times$ Faceptor [67] 4.10
FaceXFormer 88.24 FaceXFormer 72.56 FaceXFormer 4.17
Table 3: Performance comparison on facial expression recognition, face visibility prediction and age estimation. $\times$ indicates a model that doesn’t perform the task.
@@ 4.2 Implementation Details
We train our models using a distributed PyTorch setup on eight A6000 GPUs, each equipped with $48$GB of memory. The models’ backbones are initialized with ImageNet pre-trained weights and processes input images at a resolution of $224\times 224$. We employ the AdamW optimizer with a weight decay of $1e^{-5}$. All models are trained for $12$ epochs with a batch size of $48$ on each GPU, and an initial learning rate of $1e^{-4}$, which decays by a factor of $10$ at the $6^{th}$ and $10^{th}$ epochs. We train the model for three additional epochs for some tasks. For data augmentation, we randomly apply Gaussian blur, grayscale conversion, gamma correction, occlusion, horizontal flipping, and affine transformations, such as rotation, translation and scaling. The number of FaceX decoder $N$ is set to two. To ensure stable training across tasks when using multiple datasets of varying sample sizes, we equalize the representation of each task’s samples in every batch through upsampling. Additional details on our implementation are provided in the supplementary.
@@ 4.3 Main results
Methods Headpose (BIWI) Methods Landmarks (300W) Methods CelebA
Yaw Pitch Roll MAE Full Com Chal Acc.
HopeNet [74] 4.81 6.61 3.27 4.89 LAB [103] 3.49 2.98 5.19 PANDA-1 [118] 85.43
QuatNet [27] 5.49 4.01 2.94 4.15 Wing [20] 4.04 3.27 7.18 LNets+ANet [49] 87.33
FSA-Net [110] 4.27 5.49 2.93 4.14 DeCaFa [14] 3.39 2.93 5.26 SSP+SSG [30] 88.24
EVA-GCN [108] 6.01 4.78 2.98 3.98 HRNet [92] 3.32 2.87 5.15 MOON [73] 90.94
TriNet [9] 4.11 4.75 3.04 3.97 PicassoNet [100] 3.58 3.03 5.81 NSA [52] 90.61
img2pose [3] 4.56 3.54 3.24 3.78 AVS+SAN [16] 3.86 3.21 6.46 MCNN-AUX [25] 91.29
MNN [90] 3.98 4.61 2.39 3.66 LUVLi [35] 3.23 2.76 5.16 MCFA [131] 91.23
MFDNet [48] 3.40 4.68 2.77 3.62 HIH [37] 3.09 2.65 4.89 DMM-CNN [53] 91.70
TokenHPE [115] 3.95 4.51 2.71 3.72 PIPNet [29] 3.19 2.78 4.89 SSPL [80] 91.77
WHENet [124] 3.99 4.39 3.06 3.81 SLPT [106] 3.17 2.75 4.90 FaRL [123] 91.39
SwinFace [66] $\times$ $\times$ $\times$ $\times$ SwinFace [66] $\times$ $\times$ $\times$ SwinFace [66] 91.38
QFace [83] QFace [83] $\times$ $\times$ $\times$ QFace [83] 91.56
Faceptor [67] $\times$ $\times$ $\times$ $\times$ Faceptor [67] 3.16 2.75 4.84 Faceptor [67] 91.39
FaceXFormer 3.91 3.97 2.67 3.52 FaceXFormer 3.05 2.66 4.67 FaceXFormer 91.83
Table 4: Performance comparison on headpose, landmark detection, and attribute recognition. The symbol $\times$ indicates that the model does not perform the corresponding task, while – denotes that results for this dataset are not provided. Red = First Best, Blue = Second Best.
In Tab. 2, Tab. 4, Tab. 3, we present a comparative analysis of FaceXFormer with recent methods across a variety of tasks. A significant highlight of our work is its unique capability to deliver promising results across multiple tasks using a single unified model. Specifically, FaceXFormer achieves state-of-the-art performance in face parsing, with a mean F1 score of $92.01$ on CelebAMaskHQ at a resolution of $224\times 224$, which is half the input size required by other state-of-the-art methods. Furthermore, it demonstrates superior performance in head pose estimation and landmark detection, achieving a mean MAE of $3.52$ and a mean NME of $4.67$, respectively. Additionally, FaceXFormer provides a significant performance boost in attributes prediction and visibility prediction, achieving an accuracy of $91.83$% on the CelebA dataset and $72.56$% on COFW. It also performs competitively in age estimation, achieving the second-best score of $4.17$, and achieves an accuracy of $88.24$% in facial expression recognition. The results on gender estimation across different race categories is shown in Tab. 6. We present the addtional cross-dataset results in Appendix B.
Recent models such as SwinFace [66], QFace [83], and Faceptor [67] also aim to unify tasks but only address a subset of them. These models tend to exclude complex tasks such as segmentation, head pose estimation, and landmark prediction due to their conflicting training objectives. In contrast, FaceXFormer, with its learnable task-specific tokens, seamlessly unifies these tasks and achieves state-of-the-art performance across them. It outperforms previous multi-task models in tasks such as segmentation, head pose estimation, landmark prediction, attributes prediction, and visibility prediction, while achieving the second-best performance in age estimation. In this work, we simultaneously train for nine heterogeneous tasks, presenting a more formidable challenge than previous approaches. This difficulty arises primarily from the distinct and diverse nature of the tasks, which require task-specific features, pre-processing steps, and often conflicting training objectives. Despite these challenges, FaceXFormer effectively handles multiple tasks, delivering SOTA or competitive performance, thereby establishing itself as a unified SOTA model.
4.4 Qualitative “in-the-wild” results
In this section, we present the qualitative results of FaceXFormer on randomly selected “in-the-wild” images. We select three random images and showcase the results for face parsing, head pose estimation, landmarks prediction, age estimation, gender and race classification, and attributes prediction in Figure 3. Notably, the model successfully performs complex tasks such as face segmentation, head pose estimation, and landmark prediction, even when the input samples are out-of-distribution. Furthermore, FaceXFormer can be effectively used as a tool to generate multiple annotations for each image, making it valuable for various downstream tasks. These results highlight FaceXFormer’s robust performance in challenging, real-world scenarios.
5 Ablation Study
Backbone Seg Reg Cls FPS Params
MobileNet 91.21 4.64 88.22 39.76 25.32
ResNet101 91.49 4.37 88.91 34.98 65.54
ConvNext-B 92.08 4.35 89.09 36.61 110.19
Swin-B 92.01 4.12 90.03 33.21 109.29
Table 5: Effect of different backbones on performance and FPS.
In this section, we explore the impact of various backbones and their sizes on performance. Additionally, we highlight that the proposed model exhibits minimal bias compared to other models by performing age and gender prediction across different various. We provide additional ablation of different components in our model in the supplementary.
5.1 Effect of different backbone and different size.
In Table 5, we present experiments analyzing the impact of different backbones and their sizes on FaceXFormer. We group head pose estimation, landmarks prediction, and age estimation into the regression (Reg) category, while attributes prediction and facial expression recognition are categorized as classification (Cls). ConvNeXt achieves the best performance in segmentation with an F1 score of $92.08$%. The Swin Transformer backbone excels in both regression and classification tasks, with a mean error of $4.12$ and a mean accuracy of $90.03$%, respectively. In contrast, MobileNet demonstrates the lowest performance metrics, including an F1 score of $91.21$% and a mean error of $4.64$, highlighting its limitations in handling larger, more complex datasets due to its smaller receptive field compared to the Swin Transformer. The selection of the Swin Transformer as the backbone for FaceXFormer is driven by its superior scalability and global contextual understanding, both of which are essential for facial analysis tasks. Furthermore, FaceXFormer achieves real-time performance at $33.21$ FPS.
Figure 3: FaceXFormer predictions on “in-the-wild” images
5.2 Bias Analysis and Ethical Considerations
In our work, we utilize a total of $12$ datasets for training and evaluation. We obtained these datasets following the procedures stated on their respective pages and signed the license agreements if and when necessary. As we train our models on multiple datasets designed for different tasks, the subjects across different age groups, genders, and races is not equal. This imbalance may introduce bias in the model. Therefore, we provide an analysis using the FairFace [31] dataset, which is balanced in terms of age, gender and race. We follow [68] and define the ”Non-white” group to include multiple racial categories: ”Black”, ”Indian”, ”East Asian”, ”Southeast Asian”, ”Middle Eastern” and ”Latino”. As can be seen from Table 6, FaceXFormer shows the smallest performance discrepancy across different racial groups and exhibits minimal bias compared to other models despite being trained on fewer data points. This can be attributed to race estimation being the task in co-training.
@@ 5.2 Bias Analysis and Ethical Considerations
Table 6: Age and gender accuracy w.r.t race groups on FairFace
6 Discussions
Broader Impact: The world is moving towards transformers because of its potential to model large amounts of data [6, 34, 5]. Presently, the face community lacks large-scale annotated datasets to train foundational models capable of performing a wide spectrum of facial tasks. The largest clean dataset, WebFace42M [130], lacks annotations for face parsing, landmarks detection and attributes recognition. FaceXFormer can be used as an annotator for large-scale data, and can be continually improved through successive rounds of annotation and fine-tuning. We aim to propel the face community towards developing foundation models that cater to a variety of facial tasks. Additionally, FaceXFormer is a lightweight model that provides real-time output based on task-specific queries and can be appended with existing facial systems to provide additional information. It can also serve as a valuable tool in surveillance, and provide auxiliary information for subject analysis and image retrieval.
Limitations: We recognize certain limitations of the proposed FaceXFormer, particularly the requirement for complete retraining to add a new task, which may reduce its flexibility. Furthermore, although it includes token support for multiple tasks, it lacks interactivity and full promptability. Nonetheless, FaceXFormer is distinct in its ability to handle up to $9$ heterogeneous tasks within a single model, achieving state-of-the-art or competitive performance across these tasks in real-time, making it well-suited for deployment. Future work will focus on developing a large-scale pre-trained foundational model with zero-shot capabilities.
@@ 7 Conclusion
In conclusion, the FaceXformer introduces a novel end-to-end unified model that efficiently handles a comprehensive range of facial analysis tasks in real-time. By adopting a transformer-based encoder-decoder architecture and treating each task as a learnable token, our approach successfully integrates multiple tasks within a single framework. The proposed parameter-efficient decoder, FaceX, enhances the model’s ability to learn robust and generalized face representations across diverse tasks. Our comprehensive experiments demonstrate that the proposed model achieves state-of-the-art performance across multiple facial analysis tasks. Additionally, training on multiple datasets leads to better representation learning. In conclusion, we demonstrate that facial tasks can be treated as tokens, leading to the unification of tasks; following this, we hope our work provides a foundation for developing large models capable of performing multiple facial analysis tasks.
8 Acknowledgment